cable lifting an elevator

The cable lifting an elevator is wrapped around a 1.1-m-diameter cylinder that is turned by the elevator's motor. The elevator is moving upward at a speed of 2.3 m/s . It then slows to a stop, while the cylinder turns one complete revolution.


How long does it take for the elevator to stop?
Answer #1

Circumfrence of the cylinder: C = 2 \pi r

C = (2)(3.14)(1.1/2) = 3.4558 m


v2 = v02 + 2a(x)


In this case the circumfrence length will be ∆x

v = 2.3 m/s

v= 0 


2.32 = 0 + 2a(3.4558)

a = 5.29/(2*3.4558)

a = 0.76539


v = v0 + at


2.3 = 0 + (0.76539)t

t = 2.3/0.732.6

t = 3.0 s




answered by: 0bliviant
Answer #2

IMG_20211116_110023_7~2.jpg

answered by: SYED HAKEEM
Answer #3

IMG_20211116_110023_7~2.jpg

answered by: SYED HAKEEM